OpenStudy (anonymous):

Please help me! Choose the value for x that makes the open sentence true. 7 + 6 • x = 42 + x2 (Points : 1) 1.5 2.4 3.3 4.2

OpenStudy (igreen):

I think you made a mistake..

OpenStudy (igreen):

Is this it? \(7 + 6 \cdot x = 42 + x^2\)

OpenStudy (anonymous):

yes , i just checked

OpenStudy (anonymous):

two ways one: plug each number in and see which one works two: do the work 7+6x=42+x2 0=x2-6x+42-7 0=x2-6x+35

OpenStudy (anonymous):

if these are your only options GG

OpenStudy (igreen):

Oh, you have to find the zeroes of the function?